5. Dave rides the different routes on bike A and bike B. He wants to figure out if bike A is faster than bike B. Perform a hypothesis test for this using the data below. (5 points) Route Bike A (minutes) Bike B (minutes) 152 121 135 143 29 95 67 78 45 47 25 161 122 129 145 31 94 69 80 43 48 25 Mountains Flat A Canada A Canada B Michigan Work Cali A Cali B Cali C Work 1 Work 2 Daycare ride The standard deviation of differences is s-3.62. Hint: This is a one-tail test.
Answer #1
  • EXPALANTION
  • Here since the same samples of routes has been use to test the speed we will use paired samples test the hypothesis.
  • Here the random variables follow normal distribution , since the sample size is small and population standard deviation is unknown we will use t distribution to test the statistics
  • Here Let XD be the random variable denoting the speed of the bikes A and B respectively in distance per minute .
  • The differences is given by
  • bike A bike B A-B
    152 161 -9
    121 122 -1
    135 129 6
    143 145 -2
    29 31 -2
    95 94 1
    67 69 -2
    78 80 -2
    45 43 2
    47 48 -1
    25 25 0
  • The mean of the difference is given by
  • \bar{X}=\frac{\sum X}{n}=-0.90909
  • The standard deviation is given by
  • =3.62 as in question.
  • TO TEST THE HYPOTHESIS.
  • The null hypothesis is given by the that there is no difference in the speed covered by bikes A and B
  • Let the alternate hypothesis be that bike A is faster that bike B.
  • Ho:\mu d=0 vs H1:\mu d<0
  • TEST STATISTICS IS GIVEN BY
  • \frac{\bar{xd}-\mu d }{\frac{s}{\sqrt{n}}}\sim t_{n-1,\alpha }
  • \frac{-0.90909-0}{3.62/\sqrt{11}}=-0.8329
  • CONCLUSION
  • The upper and lower critical value at 5% significant level for t distribution with n-1=11-1=10 degrees of freedom \pm 1.812
  • The test statistics obatined lies between the upper and lower critical value.
  • Hence we do not have enough evidence to reject the null hypothesis.
  • Hence we conclude that there is no difference in the speed covered by bikes A and B